# FlatQuant: Flatness Matters for LLM Quantization

### v2

@@ 3.3 Efficient Kernel Design
We design an efficient kernel for FlatQuant that integrates both affine transformations and quantization into a single operation. This design is motivated by two key factors. First, $\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2}$ exhibits low computational intensity after Kronecker decomposition, making both prefill and decoding predominantly memory-bound. Second, the quantization is also known to be memory-bound.
To address these issues, we fuse $\mathcal{Q}(\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2})$ into a single kernel using OpenAI Triton (Tillet et al., 2019). Specifically, we load the entire $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ into SRAM. Each thread block slices a tiling block $\bar{\mathbf{X}}\in\mathbb{R}^{n_{1}\times n_{2}}$ from $\tilde{\mathbf{X}}$, performs the matrix multiplication $\mathbf{P}_{1}\bar{\mathbf{X}}\mathbf{P}_{2}$, and quantizes the results on the fly. Throughout this process, all intermediate results are stored in SRAM before finally being written back to the global memory. This design thereby eliminates redundant memory accesses of intermediate results and reduces the kernel launch overhead. Further details of the kernel design are provided in the Appendix B.2.
Finally, given the output above, we follow QuaRot (Ashkboos et al., 2024) to adopt the CUTLASS kernel for INT4 matrix multiplication, and FlashInfer (Ye, 2023) for KV cache quantization.
@@ 4.1 Settings
Evaluation and Baselines.
We evaluate FlatQuant on the LLaMA-2(7B/13B/70B) (Touvron et al., 2023) models and the LLaMA-3(8B/70B) (Dubey et al., 2024) models. Following previous works (Shao et al., 2023; Ashkboos et al., 2024), we report the perplexity (PPL) of language generation tasks on the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets. For commonsense reasoning tasks, we use six zero-shot evaluation tasks, including ARC-Challenge, ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), PIQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021). Additionally, we evaluate multi-turn conversation ability on LLaMA-3.1-8B-Instruct using MT-Bench, with GPT-4o as the evaluator. For baselines, we compare FlatQuant against popular INT4 post-training quantization methods, including SmoothQuant (Xiao et al., 2023), OmniQuant (Shao et al., 2023), AffineQuant (Ma et al., 2024), QUIK-4B (Ashkboos et al., 2023), and two recent state-of-the-art methods QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024b).
Implementation Details.
We implement FlatQuant based on Huggingface (Wolf, 2019) and PyTorch (Paszke et al., 2019). For optimization, we adopt the AdamW optimizer with an initial learning rate of 5e-3 and employ a cosine annealing learning rate decay schedule. Specifically, the learning rate for clipping thresholds is 5e-2. FlatQuant is trained for 15 epochs on a calibration set comprising 128 sentences from WikiText-2, each sampled with 2048 tokens. The batch size is set to 4. The default calibration procedure costs approximately 26GB of GPU memory and about 0.9 hours for LLaMA-3-8B on a single GPU. FlatQuant is robust to initialization, and we employ random affine transformation matrices as the starting point. Further details about implementation and calibration time are provided in Appendix B.1.
Quantization.
We adopt per-channel and per-token symmetric quantization for weights and activations, respectively. For KV cache quantization, we utilize group-wise asymmetric quantization with a group size of 128. This matches the head dimension of LLaMA, as suggested in previous studies (Zhao et al., 2024; Ashkboos et al., 2024), to effectively leverage the memory-bound characteristics of self-attention. By default, FlatQuant employs round-to-nearest (RTN) as the weight quantizer. For a fair comparison with QuaRot and SpinQuant, we also report weight quantization using GPTQ, which uses the same calibration data for both closed-form weight updates and training.
@@ 4.2 Main Results
We extend our evaluation to six zero-shot commonsense QA tasks, as shown in Table 2. For a fair comparison, we reproduce QuaRot  [^1] https://github.com/spcl/QuaRot and SpinQuant  [^2] https://github.com/facebookresearch/SpinQuant with their official implementations and released checkpoints, evaluating all methods with the same version of lm-eval-harness framework (Gao et al., 2021). As can be seen, FlatQuant significantly narrows the performance gap between quantized models and the FP16 baseline. Specifically, for larger models such as LLaMA-2-70B, the accuracy loss of FlatQuant is only 0.43%, which is amazing with such a low bit quantization setting. The recently released LLaMA-3 models have been shown to be more challenging for quantization (Huang et al., 2024). Nonetheless, FlatQuant continues to perform well, with an accuracy loss of 2.00% for LLaMA-3-8B and 0.94% for LLaMA-3-70B. Notably, while QuaRot with RTN largely lags behind QuaRot with GPTQ by an average accuracy gap over 4%, FlatQuant with RTN can already obtain comparable results to GPTQ.
FP16 8.17 8.10 5.05 7.00 6.10 8.67 8.50 8.91 7.60
QuaRot 7.20 6.90 3.90 5.30 4.05 6.70 6.05 7.80 5.99
FlatQuant 7.95 7.35 4.70 7.20 4.80 7.60 7.20 8.70 6.94
Results on MT-Bench.
We evaluate FlatQuant on MT-Bench with LLaMA-3.1-8B-Instruct model in Table 3. While FlatQuant trails behind the FP16 baseline in coding and STEM, it consistently outperforms QuaRot with GPTQ across all categories, narrowing the gap between the quantized model and the FP16 baseline. Notably, for math problems, FlatQuant matches the FP16 baseline’s score, exceeding QuaRot by 1.9 points. More evaluations are provided in Appendix C.2.
@@ 4.3 Inference Latency
Figure 4 shows the prefill and decoding speedup of FlatQuant across different batch sizes, with 2048 and 256 tokens for prefill and decoding, respectively. It can be found that even without kernel fusion, FlatQuant acheives comparable speed-up with QuaRot, thanks to kronecker decomposition of affine transformations. With kernel fusion, FlatQuant can achieve up to 2.30x speedup for prefill and 1.76x speedup for decoding under the batch size of 64, which is apparently faster than QuaRot (Ashkboos et al., 2024). Although there is still a minor gap compared to the vanilla INT4 quantization, it significantly enhances accuracy and facilitates the deployment of INT4 LLMs in real-world applications.
Kronecker Decomposition: Sizes and Perplexities.
In Figure 6, we examine the impact of different decomposed matrix sizes in Equation 3 on model performance and speedup. As shown, the varying sizes of Kronecker decomposition significantly affect speedup, but have limited impact on the perplexity of generated text. The speedup peaks when $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are of equal size (i.e., $n_{1}=n_{2}=\sqrt{n}=64$), as predicted by our theoretical analysis in Section 3.1. When $n_{2}$ exceeds 64, the speedup quickly decreases due to irregular memory access patterns for activations. These results further demonstrate FlatQuant’s effectiveness in minimizing inference overhead while maintaining quantization accuracy through matrix decomposition.
@@
We now investigate the impact of the five online transformations (i.e., $\{\mathbf{P}_{a},\mathbf{P}_{o},\mathbf{P}_{h},\mathbf{P}_{ug},\mathbf{P}_{d}\}$) in FlatQuant on the overall speedup, as shown in Figure 6. Even with five per-layer transformations, FlatQuant results in a minimal 0.07x end-to-end slowdown, significantly outperforming QuaRot’s 0.26x with just three Hadamard transformations. Specifically, FlatQuant’s $\mathbf{P}_{d}$ causes a 0.04x slowdown due to large FFN intermediate sizes, compared with QuaRot’s 0.17x. Meanwhile, $\mathbf{P}_{o}$ results in a 0.01x slowdown, versus QuaRot’s 0.1x. The rest transformations (i.e., $\mathbf{P}_{a}$ and $\mathbf{P}_{ug}$) have an insignificant impact of less than 0.01x. Finally, it can be found that even without kernel fusion, the additional transformations in FlatQuant is still on par with QuaRot, thanks to Kronecker decomposition.
@@ 4.4 Discussions
LT PS LCT WikiText-2 C4 ARC-C ARC-E HellaSwag LAMBADA PIQA Winogrande Avg
1266.60 936.41 25.26 28.62 27.04 1.26 51.80 51.93 30.99
8.50 13.51 44.97 71.38 73.17 67.05 76.88 67.48 66.82
7.95 12.74 44.20 71.89 74.21 68.72 77.15 66.30 67.08
7.11 11.47 49.32 76.14 76.30 72.17 78.89 71.51 70.72
6.98 11.13 50.00 75.80 76.80 72.91 79.16 72.69 71.23
@@
We conduct ablation studies for FlatQuant focusing on its main components: 1) learnable transformation (LT); 2) per-channel scaling (PS); and 3) learnable clipping thresholds (LCT). Starting from RTN as a baseline, we evaluate the impact of each component on perplexity and the average accuracy on zero-shot QA tasks, with LLaMA-3-8B model. As shown in Table 4, enabling LT significantly enhances the accuracy of the quantized model, reducing PPL from 1266.60 to 8.50 on WikiText-2. This shows LT is capable of adaptively flattening the distribution of weight and activation values. Additionally, incorporating PS and LCT further improves PPL by 0.55 and 0.84, respectively, demonstrating the necessity of each component to refine the model performance.
LLaMA-3-8B WikiText-2 PPL C4 PPL
W4A16 W3A16 W4A16 W3A16
FP16 6.14 9.45
RTN 8.70 2.2E3 14.00 5.6E3
GPTQ 7.00 13.00 11.80 45.90
GPTQ-g128 6.50 8.20 10.40 13.70
AWQ 7.10 12.80 10.10 16.80
QuIP 6.50 7.50 11.10 11.30
FlatQuant-RTN 6.54 7.78 10.17 12.64
FlatQuant-GPTQ 6.48 7.52 10.28 12.91
Table 5: Weight-only quantization results on LLaMA-3-8B model.
W4 A4 KV4 WikiText-2 PPL C4 PPL QA Acc
6.14 9.45 73.23
6.56 10.25 72.92
6.49 10.13 72.20
6.23 9.61 73.43
6.98 11.13 71.23
Other Quantization Schemes.
Although the main results above focus mostly on weight-activation quantization, FlatQuant can be easily applied to other quantization schemes. The results of weight-only quantization against several state-of-the-art baselines are presented in Table 6. FlatQuant again obtains leading accuracy compared with leading baselines. For additional results on KV cache quantization and extreme low-bit quantization, please refer to Appendix C.2.
Train One and Get More.
We demonstrate that the affine transformations learned from weight-activation quantization can be directly applied to other quantization settings, such as weight-only or KV cache quantization, with surprisingly strong performance. The associated results are presented in Table 6. For instance, the results labeled as “W4” are comparable to those in Table 6 that are specifically trained for weight-only quantization. This significantly saves time when applying FlatQuant to different quantization settings, as only one set of transformation matrices is saved.
Due to space constraints, we provide additional discussions such as the impact of calibration data, and the effect of learnable clipping in Appendix C.3. More visualizations on the flatness of transformed weights and activations, and the quantization error landscapes are in Appendix D.
@@ 5 Conclusions
In this study, we revisit the importance of flat weights and activations for effective quantization, and find existing solutions still produce steep outspread values after the pre-quantization transformation. Therefore, we introduce FlatQuant, a novel post-training quantization method with the purpose of identifying fast and learnable transformations for each linear layer, to promote the flatness of weights and activations. Extensive experiments demonstrate the superiority of FlatQuant, e.g., with less than 1% accuracy drop for W4A4 quantization on the LLaMA-3-70B. Our efficient kernel fusion integrates the affine transformation and quantization, reducing the transformation overhead and bringing up to 2.3x and 1.7x speedup over FP16 inference at the prefill and decoding stages, respectively. We hope this work advances the practical application of low-bit quantization for LLMs.
@@ Appendix A Related Work
Quantization is a crucial technique for reducing memory footprint and accelerating inference by employing fewer bits for storage and computation, especially for the application of LLMs. Unlike previous models, LLMs are shown to exhibit outliers in activation and massive outliers in pivot tokens (Wei et al., 2022; Dettmers et al., 2022; Liu et al., 2024a; Sun et al., 2024), which can severely degrade quantization accuracy. To eliminate the negative impact of outliers, pre-quantization transformations have been widely adopted in weight-activation quantization (Xiao et al., 2023; Wei et al., 2023; Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b) as well as in fully quantized training (Xi et al., 2023). Additionally, several weight-only quantization methods (Lin et al., 2023; Chee et al., 2024; Tseng et al., 2024) incorporate pre-quantization transformations. Searched or learnable clipping thresholds for weights or activations (Lin et al., 2023; Duanmu et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b; Shao et al., 2023) are also explored to eliminate outliers.
@@
Recent research (Xi et al., 2023; Tseng et al., 2024; Ashkboos et al., 2024) has shown that the Hadamard transformation is effective in eliminating outliers and lowering quantization error by redistributing outliers across all channels through matrix multiplication. QuaRot (Ashkboos et al., 2024) is the first to apply Hadamard transformation in the LLM W4A4 PTQ setting, while SpinQuant (Liu et al., 2024b) exploits learnable orthogonal matrices with model-level loss to further alleviate outliers.
@@
Inspired by SmoothQuant, AWQ (Lin et al., 2023) introduces activation-aware per-channel scaling to reduce quantization errors in weight-only quantization. QUIP (Chee et al., 2024) and its extension, QUIP# (Tseng et al., 2024), leverage random rotation matrices or Hadamard transformations to enhance incoherence in weight-only quantization. In fully quantized training task, Xi et al. (2023) propose to utilize a block-diagonal transformation consisting of Hadamard matrices to reduce the quantization error.

### v3

@@ 3.3 Efficient Kernel Design
We design an efficient kernel for FlatQuant that integrates both affine transformations and quantization into a single operation. This design is motivated by two key factors. First, $\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2}$ exhibits low computational intensity with Kronecker product of two lightweight matrices, making both prefilling and decoding predominantly memory-bound. Second, the quantization is also known to be memory-bound.
To address these issues, we fuse $\mathcal{Q}(\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2})$ into a single kernel using OpenAI Triton (Tillet et al., 2019). Specifically, we load the entire $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ into SRAM. Each thread block slices a tiling block $\bar{\mathbf{X}}\in\mathbb{R}^{n_{1}\times n_{2}}$ from $\tilde{\mathbf{X}}$, performs the matrix multiplication $\mathbf{P}_{1}\bar{\mathbf{X}}\mathbf{P}_{2}$, and quantizes the results on the fly. Throughout this process, all intermediate results are stored in SRAM before finally being written back to the global memory. This design thereby eliminates redundant memory accesses of intermediate results and reduces the kernel launch overhead. Finally, given the output above, we follow QuaRot (Ashkboos et al., 2024) to adopt the CUTLASS kernel for INT4 matrix multiplication, and FlashInfer (Ye, 2023) for KV cache quantization. Further details of the kernel design are provided in the Appendix B.3.
@@ 4.1 Settings
Evaluation and Baselines.
We primarily evaluate FlatQuant on the series of LLaMA-2 (Touvron et al., 2023) and LLaMA-3 (Dubey et al., 2024) models, and results on Qwen and DeepSeek families can be found in Appendix C.1. Following previous works (Shao et al., 2023; Ashkboos et al., 2024), we report the perplexity (PPL) of language generation tasks on the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets. For commonsense reasoning tasks, we use six zero-shot evaluation tasks, including ARC-Challenge, ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), PIQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021). We compare FlatQuant against popular INT4 post-training quantization methods, including SmoothQuant (Xiao et al., 2023), OmniQuant (Shao et al., 2023), AffineQuant (Ma et al., 2024), QUIK-4B (Ashkboos et al., 2023), and two recent state-of-the-art methods QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024c).
Implementation Details.
We implement FlatQuant based on Huggingface (Wolf, 2019) and PyTorch (Paszke et al., 2019). We adopt the AdamW optimizer with an initial learning rate of 5e-3 and employ a cosine annealing learning rate decay schedule. The learning rate for clipping thresholds is 5e-2. FlatQuant is trained for 15 epochs on a calibration set comprising 128 sentences from WikiText-2, each sampled with 2048 tokens. The batch size is set to 4. The default calibration procedure costs approximately 26GB of GPU memory and about 0.9 hours for LLaMA-3-8B on a single GPU. FlatQuant is robust to initialization, and we employ random affine transformation matrices as the starting point. Further details about implementation and calibration time are provided in Appendix B.1.
Quantization.
We adopt per-channel and per-token symmetric quantization for weights and activations, respectively. For fair comparisons with QuaRot and SpinQuant, we employ both round-to-nearest (RTN) and GPTQ as the weight quantizer, where GPTQ uses the same calibration data for both closed-form weight updates and training. Nonetheless, our empirical results suggest that FlatQuant with RTN is sufficient to be competitive. FlatQuant can be also used for KV cache quantization, where group-wise asymmetric quantization with the size of 128 is applied. This matches the head dimension of LLaMA, as suggested in previous studies (Zhao et al., 2024; Ashkboos et al., 2024), to leverage the memory-bound characteristics of self-attention.
@@ 4.2 Main Results
We extend our evaluation to six zero-shot commonsense QA tasks, as shown in Table 2. For a fair comparison, we reproduce QuaRot  [^1] https://github.com/spcl/QuaRot and SpinQuant  [^2] https://github.com/facebookresearch/SpinQuant with their official implementations and released checkpoints, evaluating all methods with the same version of lm-eval-harness framework (Gao et al., 2021). As can be seen, FlatQuant significantly narrows the performance gap between quantized models and the FP16 baseline. Specifically, while the LLaMA-3 models are shown to be challenging for quantization (Huang et al., 2024), FlatQuant perform well with an accuracy loss of 2.00% for LLaMA-3-8B and 0.94% for LLaMA-3-70B. Notably, while QuaRot with RTN largely lags behind QuaRot with GPTQ by an average accuracy gap over 4%, FlatQuant with RTN can already obtain comparable results to GPTQ.
Due to limited space, we leave more experimental results in Appendix C, such as exploration to other quantization settings in Appendix C.3, results on more LLM architectures in Appendix C.1, and MT-bench evaluations in Appendix C.2.
@@ 4.3 Inference Latency
Figure 4 shows the prefill and decoding speedup of FlatQuant across different batch sizes, with 2048 and 256 tokens for prefill and decoding, respectively. It can be found that even without kernel fusion, FlatQuant acheives comparable speed-up with QuaRot, thanks to the Kronecker product of two lightweight matrices. With kernel fusion, FlatQuant can achieve up to 2.30x speedup for prefill and 1.76x speedup for decoding under the batch size of 64, which is apparently faster than QuaRot (Ashkboos et al., 2024). Although there is still a minor gap compared to the vanilla INT4 quantization, it significantly enhances accuracy and facilitates the deployment of INT4 LLMs in real-world applications.
Kronecker Product: Sizes and Perplexities.
In Figure 6, we examine the impact of different decomposed matrix sizes in Equation 3 on model performance and speedup. As shown, the varying sizes of matrices for the Kronecker product significantly affect speedup. However, they have limited impact on the perplexity of generated text. The speedup peaks when $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are of equal size (i.e., $n_{1}=n_{2}=\sqrt{n}=64$), as predicted by our theoretical analysis in Section 3.1. When $n_{2}$ exceeds 64, the speedup quickly decreases due to irregular memory access patterns for activations. These results further demonstrate FlatQuant’s effectiveness in minimizing inference overhead while maintaining quantization accuracy with the Kronecker product.
@@
We now investigate the impact of the five online transformations (i.e., $\{\mathbf{P}_{a},\mathbf{P}_{o},\mathbf{P}_{h},\mathbf{P}_{ug},\mathbf{P}_{d}\}$) in FlatQuant on the overall speedup, as shown in Figure 6. Even with five per-layer transformations, FlatQuant results in a minimal 0.07x end-to-end slowdown, significantly outperforming QuaRot’s 0.26x with just three Hadamard transformations. Specifically, FlatQuant’s $\mathbf{P}_{d}$ causes a 0.04x slowdown due to large FFN intermediate sizes, compared with QuaRot’s 0.17x. Meanwhile, $\mathbf{P}_{o}$ results in a 0.01x slowdown, versus QuaRot’s 0.1x. The rest transformations (i.e., $\mathbf{P}_{a}$ and $\mathbf{P}_{ug}$) have an insignificant impact of less than 0.01x. Finally, it can be found that even without kernel fusion, the additional transformations in FlatQuant is still on par with QuaRot, thanks to the Kronecker product of two lightweight matrices.
Figure 7: Flatness and mean squared quantization error (MSE) of different Transformer blocks in LLaMA-3-8B during FlatQuant’s training process. The metric of flatness is calculated as the sum of Euclidean distances $\|\mathbf{d}-\mathbf{d}^{{}^{\prime}}\|_{2}$ for all weights and activations within a Transformer block.
@@ 4.4 Discussions
LT PS LCT WikiText-2 C4 Avg
1266.60 936.41 30.99
8.50 13.51 66.82
7.95 12.74 67.08
7.11 11.47 70.72
6.98 11.13 71.23
@@
We conduct ablation studies for FlatQuant focusing on its main components: 1) learnable transformation (LT); 2) per-channel scaling (PS); and 3) learnable clipping thresholds (LCT). Starting from RTN as a baseline, we evaluate the impact of each component on perplexity and the average accuracy on zero-shot QA tasks based on LLaMA-3-8B. As shown in Table 3, enabling LT significantly enhances the accuracy of the quantized model, reducing PPL from 1266.60 to 8.50 on WikiText-2. This shows LT is capable of adaptively flattening the distribution of weights and activations. Additionally, incorporating PS and LCT further improves PPL by 0.55 and 0.84, respectively, demonstrating the necessity of each component to enhance the performance. Due to space limitation, we leave the more comprehensive ablation in Appendix C.4.
FlatQuant Leads to Flatness.
To further analyze how FlatQuant promotes flatness, we quantitatively evaluate the flatness of weights and activations by analyzing their channel-wise magnitude distributions. Specifically, each distribution is represented as a one-dimensional vector $\mathbf{d}$, as illustrated in Figure 1. We measure flatness with the mean squared error (MSE) between the observed distribution $\mathbf{d}$ and an idealized perfectly flat distribution $\mathbf{d}^{\prime}$. The flat distribution $\mathbf{d}^{\prime}$ is defined such that all channels possess equal magnitudes and the same $\ell_{2}$ norm as $\mathbf{d}$, i.e., $\mathbf{d}^{\prime}=\frac{\|\mathbf{d}\|_{2}}{\sqrt{N}}\cdot\mathbf{1}_{N}$, where $N$ is the number of channels and $\mathbf{1}_{N}$ is an $N$-dimensional vector with all entries equal to one. The Euclidean distance $\|\mathbf{d}-\mathbf{d}^{{}^{\prime}}\|_{2}$ thus serves as our flatness metric, where smaller values indicate distributions closer to uniformity across all channels. In Figure 7, we visualize the evolution of flatness and the training objective (Equation 4) across different Transformer blocks of LLaMA-3-8B during training. With the decreasing training loss, the channel distributions become increasingly flat. This indicates that FlatQuant learns better transformations to obtain a flatter distribution which ultimately contributes to smaller quantization error, i.e., flatness matters for LLM quantization.
Due to space constraints, we provide additional experiments and discussions in Appendix C.5, including the impact of calibration data, the effect of learnable clipping and mixed-precision schemes. We also analyze inference memory consumption in Appendix C.7. Additional visualizations of flatness and quantization error landscapes are provided in Appendix D.1 and D.2, respectively.
@@ 5 Conclusions
In this study, we revisit the importance of flat weights and activations for effective quantization, and find existing solutions still produce steep outspread values after the pre-quantization transformation. Therefore, we introduce FlatQuant, a novel post-training quantization method with the purpose of identifying fast and learnable transformations for each linear layer, to promote the flatness of weights and activations. Extensive experiments demonstrate the superiority of FlatQuant, e.g., with less than 1% accuracy drop for W4A4 quantization on the LLaMA-3-70B. Our efficient kernel fusion integrates the affine transformation and quantization, bringing up to 2.3x and 1.7x speedup over FP16 inference at the prefill and decoding stages, respectively. We hope this work advances the practical application of low-bit quantization for LLMs.
Impact Statement
This paper presents work aimed at advancing the field of Machine Learning by improving the efficiency of LLMs through enhanced quantization techniques. These improvements have the potential to lower operational costs and energy consumption, enabling broader access to AI, fostering more equitable use of advanced technologies. However, it is important to recognize that quantization does not address inherent societal biases in training data. Careful consideration and responsible deployment of quantized models are necessary to mitigate ethical risks and ensure their fair use.
@@ Appendix A Related Work
Quantization is a crucial technique for reducing memory footprint and accelerating inference by employing fewer bits for storage and computation. Unlike pruning (Ma et al., 2023; Sun et al., 2023; Zhang et al., 2024; Chen et al., 2025), quantization does not alter the network architecture and is usually more competitive in performance under the same compression ratio. Different from the previous pretrained language models (Shen et al., 2020; Bai et al., 2020, 2022), LLMs are shown to exhibit outliers in activation and massive outliers in pivot tokens (Wei et al., 2022; Dettmers et al., 2022; Liu et al., 2024a; Sun et al., 2024), which can severely degrade quantization accuracy, especially on on complex reasoning tasks (Liu et al., 2025). To eliminate the negative impact of outliers, pre-quantization transformations have been widely adopted in weight-activation quantization (Xiao et al., 2023; Wei et al., 2023; Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024c) as well as in fully quantized training (Xi et al., 2023). Additionally, several weight-only quantization methods (Lin et al., 2023; Chee et al., 2024; Tseng et al., 2024) incorporate pre-quantization transformations. Searched or learnable clipping thresholds for weights or activations (Lin et al., 2023; Duanmu et al., 2024; Ashkboos et al., 2024; Liu et al., 2024c; Shao et al., 2023) are also explored to eliminate outliers.
@@
Recent research (Xi et al., 2023; Tseng et al., 2024; Ashkboos et al., 2024) has shown that the Hadamard transformation is effective in eliminating outliers and lowering the quantization error by redistributing outliers across all channels through matrix multiplication. QuaRot (Ashkboos et al., 2024) is the first to apply Hadamard transformation in the LLM W4A4 PTQ setting, while SpinQuant (Liu et al., 2024c) exploits learnable orthogonal matrices with model-level loss to further alleviate outliers.
@@
Inspired by SmoothQuant, AWQ (Lin et al., 2023) introduces activation-aware per-channel scaling to reduce quantization errors in weight-only quantization. QUIP (Chee et al., 2024) and its extension, QUIP# (Tseng et al., 2024), leverage random rotation matrices or Hadamard transformations to enhance incoherence in weight-only quantization. In fully quantized training task, (Xi et al., 2023) propose to utilize a block-diagonal transformation consisting of Hadamard matrices to reduce the quantization error.